# Quasicrystalline 30 twisted bilayer graphene: Fractal patterns and electronic localization properties

### v1

@@ I Introduction
The electronic properties of bilayer systems when twisted at specific angles, have been a subject of intense research interest. In particular, twisted bilayer graphene (TBG) near to the so-called magic angles show a plethora of phenomena, such as superconductivity [1, 2, 3], magnetism [4, 5, 6, 7], and strong electronic correlation effects [8, 3]. Recently, the discovery of dodecagonal quasicrystals [9, 10] in $30^{\circ}$-twisted bilayer graphene ($30^{\circ}$-TBG) has further attracted considerable attention to TBGs, as it offers a unique platform to study the effects arising from long-range order without translational symmetry. Unlike conventional quasicrystals, where all atoms are intrinsically located within a quasiperiodic order [11, 12], a $30^{\circ}$-TBG is viewed as an extrinsic quasicrystal because its quasiperiodicity arises from the interlayer coupling between two graphene monolayers [13].
Dodecagonal $30^{\circ}$-TBG quasicrystals have been successfully synthesized on various substrates, including SiC [10, 14, 15], Pt [9], Cu [16, 17, 18, 19], and a Cu-Ni alloy [20], using methods such as chemical vapor deposition and carbon segregation. The $30^{\circ}$-TBGs are experimentally identified by low-energy electron diffraction (LEED) [10, 9, 14, 20], high-energy positron diffraction (HEPD)[15], transmission electron microscopy (TEM) [10, 18], scanning tunneling microscopy (STM) [16], and Raman spectroscopy [9, 17, 19]. Furthermore, magnetotransport measurements have confirmed the low-energy Dirac fermion nature of these systems [16, 17].
ARPES measurements reveal multiple Dirac cones with 12-fold rotational symmetry in $30^{\circ}$-TBGs [9, 10]. The emergence of mirrored Dirac cones within the Brillouin zone of each graphene layer and a bandgap at the zone boundary, attributed to generalized Umklapp scattering between the two layers, characterizes their coupling [9, 10]. STS measurements [16] reveal a suppression of density of state (DOS) consistent with the bandgap observed in previous ARPES studies, suggesting the same origin [16]. Additionally, time- and angle-resolved photoemission spectroscopy measurements show an unbalanced electron distribution in the replica Dirac cone bands [21].
The dodecagonal $30^{\circ}$-TBG quasicrystal is a direct realization of the Stampfli construction process [22], which involves equilataral triangles, squares, and rhombuses as fundamental units arranged in a dodecagonal pattern. False-colored TEM images of $30^{\circ}$-TBGs [10, 18] clearly depict the Stampfli tiles in various orientations.
@@
Recent theoretical predictions for $30^{\circ}$-TBG systems are rich and diverse. Numerical simulations suggest that the perfect superlubricity characterized by a scale-invariant sliding force related to the contact area, can be achieved for geometric sequences of dodecagonal quasicrystal tiling elements [28]. The system’ quasiperiodicity and weak interlayer coupling can originate quantum oscillations with spiral Fermi surfaces [29]. We note that the nature of interlayer hybridization selection rules governing interlayer coupling have been thoroughly investigated, via symmetry and group representation theory [30]. Unlike the only allowed equivalent hybridization in $D_{6h}$ untwisted graphene bilayer, $D_{6}$ twisted graphene bilayer permits equivalent and mixed hybridizations. Furthermore, $D_{6d}$ dodecagonal quasicrystal in $30^{\circ}$-TBG allows for both equivalent and nonequivalent hybridizations [30]. Additionally, the presence of an exchange field and a Rashba spin-orbit coupling have been proposed to induce a nontrivial topological phase in $30^{\circ}$-TBG systems [31].
In this study, we reexamine the electronic properties of $30^{\circ}$-TBGs, with focus on the spacial distribution of the localized states, using the Haydock-Heine-Kelly (HHK) recursion technique [32, 33, 34, 35]. This method relies on the nearsightedness of electronic matter (see, for instance, Ref. [36]) and, as a real-space approach, unlike the standard techniques [37], eliminates the need for periodic boundary conditions, making it ideal for analyzing strongly disordered, amorphous, and quasicrystalline electronic systems. We demonstrate that the HHK-method outperforms the previously proposed approaches for $30^{\circ}$-TBG systems.
@@ II.1 Lattice structure
Figure 1a illustrates the first Stampfli construction method obtained from the grid method [22, 10]. The quasicrystal lattice is composed of equilateral triangles, squares and rhombuses (with $30^{\circ}$ acute angle), arranged with different orientations to mantain a 12-fold rotational symmetry pattern, while lacking translational symmetry. The inflation rule of Stampfli tiles involved rotating the lattice by $15^{\circ}$ around its perfect rotational symmetry center and scaling side lengths by a factor $\sqrt{2+\sqrt{3}}$ [22]. This process generates a new lattice whose vertices coincide with the ones of the original lattice.
@@
Figure 1b presents the second construction method as a compound tessellation [28, 22]. The dodecagonal compound tessellation is decomposed into equilateral triangles and squares, as depicted in Fig. 1c. The center of each dodecagon corresponds to a vertex of a larger, similar dodecagonal element, with side lengths scaled by a factor $2+\sqrt{3}$ [22].
In the next section, we use the properties of the Stampfli construction methods to discuss electronic emergent localization features of $30^{\circ}$-TBG systems. More specifically, we employ the third inflation of Stampfli tiles, seed multiplied by a factor $\left(2+\sqrt{3}\right)^{3/2}$, as used in the first construction method. This ensures that the internal structures of each tile kind are similar [10]. Figure 2 highlights five distinct radial distances denoted as $r_{1}$, $r_{2}$, $r_{3}$, $c_{1}$ and $c_{2}$, where the Stampfli tile vertices serve as central points of emergent geometric site concentration.
Figure 2: Geometric model: The five different highlighted-color, marked as $r_{1}$, $r_{2}$, $r_{3}$, $c_{1}$, $c_{2}$, indicate an enhanced localization of carbon atom sites at five different radial distances, using as localization references (a) the Stampfli tile vertices and (b) the dodecagonal centers obtained by the first and second construction methods, respectively. Red and blue lines correspond to the top and bottom graphene layers, respectively.
Central points formed by a single kind of Stampli tiles, coinciding with the main twist axis perpendicular to the plane, exhibit a perfect 12-fold symmetry. In contrast, central points formed by the vertices of different Stampfli tiles, corresponding to carbon atom sites at the same radial distances, are characterized by imperfect rings. Most imperfect rings exhibit an imbalance in emergent state measures, as detailed in Sec. III. The imbalance primarily originates from the distinct symmetry groups of the Stampfli tiles: rhombuses, triangles and squares posses twofold, threefold and fourfold symmetry, respectively, see Fig. 2a. Near-perfect rings are readily identified by the dodecagonal centers used in the second construction method, red circles in Fig. 1b, where center point regions exhibit pronounced 12-fold symmetry and balanced emergent state measures.
@@ II.2 Model Hamiltonian
$V_{pp\pi}(R)=$ $\,V^{0}_{pp\pi}\exp\left(-\frac{R-a_{0}}{\delta}\right)\quad\mbox{and}$
@@ II.3 Numerical method
We compute the local spectral function of $30^{\circ}$-TBG quasicristals using the Haydock-Heine-Kelly (HHK) recursion technique [32, 33, 34, 35]. The HHK method is an $O(\mathcal{N})$ approach which consists of a recursive procedure [34] that transforms an arbitrary sparse Hamiltonian matrix in a tridiagonal form. Next, one evaluates the diagonal Green’s function, $G_{ii}^{r}(\epsilon)\equiv G^{r}(\textbf{R}_{i},\textbf{R}_{i},\epsilon)$, by a continued fraction expansion, that is very amenable for numerical calculation. A detailed presentation of the method and its implementation can be found, for instance, in Refs. [34, 35].
@@ III Results
Using the HHK method we confirm that the equivalent sites at identical radial distance share the same local spectral function due to the most relevant symmetry operators, as discussed in Sec. II.1. Consequently, our calculations require the computation of only one twenty-forth of the total number of carbon atoms, within a specified circular region of the $30^{\circ}$-TBG dodecagonal quasicrystal.
@@ III.1 Local electronic structures
In Fig. 4 error bars quantify the standard deviation from the mean value, primarily due to the peaks associated to the valence band singularities that vary along the radial distance. For a given kind of Stampfli tile, the first inflation shows fluctuations of order $10^{-3}$ and $10^{-4}$ along the radial distance. By the third inflation, standard deviations are negligible for all spectral functions, regardless of radial distances. A small standard deviation indicates greater similarity with the first spectral function. Thus, the third inflation is the smallest order at which the spectral functions of the Stampfli unit structures exhibit typical patterns characteristic of the dodecagonal quasicrystal. By accounting for the normalization factor that relates the spectral functions of the different Stampfli unit structures, we observe a excellent agreement within the numerical precision. Our findings suggest that the electronic structure of the system can be calculated using only a single Stampfli unit structure.
@@ III.2 Emergent localization
Figures 6a to 6e depict real-space images of emergent localized state of a $30^{\circ}$-TBG system, corresponding to the peaks $\alpha$, $\beta$, $\gamma$, $\tau$, and $\mu$ of DOS singularities in the valence band, respectively. The sizes of the red (blue) circles are proportional to the LDOS at the top (bottom) layer site at which the circle is centered. The peak energies $\alpha$, $\beta$, $\gamma$, $\tau$ and $\mu$ are approximately $-1.163$ eV, $-0.949$ eV, $-0.702$ eV, $-0.842$ eV and $-0.44$ eV, respectively. All emergent states exhibit a non-periodic dodecagonal pattern and can be spatially represented by a quasicrystal lattice model at the third inflation of the Spampfli tiles.
@@
Figures 6f to 6j present histograms illustrating the spatial distribution of states corresponding to the energy peaks $\epsilon_{i}$ with $i=\{\alpha,\beta,\gamma,\tau,\mu\}$, respectively. We analyze the variation of emergent localized states along the spacial radial distances, $d=\{r_{1},r_{2},r_{3},c_{1},c_{2}\}$, as indicated in Figure 2, where the vertices of the Stampfli tiles serve as central points. Each histogram shows the average LDOS for the energy peak $\epsilon_{i}$, $\rho(\epsilon_{i},d)=\sum_{j\in d}\rho_{j}(\epsilon_{i})$, where $j$ corresponds to the atomic sites located at a distance $d$. The average LDOS is normalized by its maximum value, $\rho_{\rm max}(\epsilon_{i},d)$.
@@
The black circles in the insets of Figs. 6a to 6c and in Figs. 6d and 6e correspond to a zoom-in view of the structures of the central point regions formed by exclusively by Stampfli rhombuses associated with the energy peaks $\epsilon_{i}$ with $i=\{\alpha,\beta,\gamma,\tau,\mu\}$, respectively. The distribution of the LDOS in each peak, $\rho_{j}(\epsilon_{i})$, ehxibits a perfect 12-fold symmetry. In the insets of Figs. 6a to 6c, the red and green circles identify regions corresponding to dodecagonal centers used in the second construction method and the vertices of different Stampfli tiles used in the first construction method, respectively. While the red circles do not locally preserve 12-fold symmetry, their distribution is close to the latter. However, the green circles deviate significantly form 12-fold symmetry, displaying an imbalance due to the different symmetry groups of the Stampfli tiles.
@@
We observe that the average LDOS at specific radial distances, $\rho(\epsilon_{i},d)$, in regions with strong imbalances resemble those in in the regions with perfect 12-fold symmetry.
@@
Figures 6a and 6f demonstrate that the $\alpha$-peak states are highly localized on the second ring of carbon atoms closest to the center points, at a radial distance $r_{2}$, with $\rho_{\rm max}(\alpha,r_{2})=0.203$ a.u. Interestingly, these second-ring atomic sites can be considered as opposite of dimer sites, due to their significant spacing.
@@
Figures 6b and 6g reveal a quasi-homogeneous emergent states distribution for the $\beta$-peak, with $\rho_{\rm max}(\beta,c_{2})=0.119$ a.u., indicating less less localization at radial distance $c_{1}$. Figures 6i and 6j show that the peaks $\tau$ and $\mu$ peaks also exhibit quasi-homogeneous emergent states distributions, with $\rho_{\rm max}(\tau,r_{1})=0.089$ a.u. and $\rho_{\rm max}(\mu,c_{1})=0.062$ a.u., respectively. While the $\tau$-peak states are less localized at $c_{1}$, the $\mu$-peak states are less localized at $r_{2}$ and $c_{2}$.
@@
Figure 6h illustrates that the $\gamma$-peak states are highly localized at a radial distance $c_{1}$, coinciding with the dimer sites, with $\rho_{\rm max}(\gamma,c_{1})=0.102$ a.u. In Figure 6c, the largest localizations are indicated by the purple disks, which correspond to overlapping red and blue disks representing the top and bottom layers, respectively.

### v2

@@ I Introduction
The electronic properties of bilayer systems when twisted at specific angles, have been a subject of intense research interest. In particular, twisted bilayer graphene (TBG) near to the so-called magic angles show a plethora of phenomena, such as superconductivity [1, 2, 3], magnetism [4, 5, 6, 7], and strong electronic correlation effects [8, 3]. Some of these interesting phenomena are also present in twisted multilayer graphene [9, 10]. Recently, the discovery of dodecagonal quasicrystals [11, 12] in $30^{\circ}$-twisted bilayer graphene ($30^{\circ}$-TBG) has further attracted considerable attention to TBGs, as it offers a unique platform to study the effects arising from long-range order without translational symmetry. Unlike conventional quasicrystals, where all atoms are intrinsically located within a quasiperiodic order [13, 14], a $30^{\circ}$-TBG is viewed as an extrinsic quasicrystal because its quasiperiodicity arises from the interlayer coupling between two graphene monolayers [15].
Dodecagonal $30^{\circ}$-TBG quasicrystals have been successfully synthesized on various substrates, including SiC [12, 16, 17], Pt [11], Cu [18, 19, 20, 21], and a Cu-Ni alloy [22], using methods such as chemical vapor deposition and carbon segregation. The $30^{\circ}$-TBGs are experimentally identified by low-energy electron diffraction (LEED) [12, 11, 16, 22], high-energy positron diffraction (HEPD)[17], transmission electron microscopy (TEM) [12, 20], scanning tunneling microscopy (STM) [18], and Raman spectroscopy [11, 19, 21]. Furthermore, magnetotransport measurements have confirmed the low-energy Dirac fermionic nature of these systems [18, 19].
Various techniques have been employed to characterize dodecagonal quasicrystalline properties of $30^{\circ}$-TBGs. Angle-resolved photoemission spectroscopy (ARPES) measurements have revealed multiple Dirac cones with 12-fold rotational symmetry in $30^{\circ}$-TBGs [11, 12], a feature that is highly sensitive to deviations from the $30^{\circ}$ twist angle [12]. The emergence of mirrored Dirac cones within the Brillouin zone of each graphene layer and a bandgap at the zone boundary, attributed to generalized Umklapp scattering between the two layers, characterizes their coupling [11, 12]. STS measurements [18] reveal a suppression of density of state (DOS) consistent with the bandgap observed in previous ARPES studies, suggesting the same origin [18]. Additionally, time- and angle-resolved photoemission spectroscopy measurements show an unbalanced electron distribution in the replica Dirac cone bands [23].
The dodecagonal $30^{\circ}$-TBG quasicrystal is a direct realization of the Stampfli construction process [24], which involves equilateral triangles, squares, and rhombuses as fundamental units arranged in a dodecagonal pattern. False-colored TEM images of $30^{\circ}$-TBGs [12, 20] clearly depict the Stampfli tiles in various orientations.
@@
Recent theoretical predictions for $30^{\circ}$-TBG systems are rich and diverse. Numerical simulations suggest that the perfect superlubricity characterized by a scale-invariant sliding force related to the contact area, can be achieved for geometric sequences of dodecagonal quasicrystal tiling elements [30]. The system’s quasiperiodicity and weak interlayer coupling can originate quantum oscillations with spiral Fermi surfaces [31]. We note that the nature of interlayer hybridization selection rules governing interlayer coupling has been thoroughly investigated, via symmetry and group representation theory [32]. Unlike the only allowed equivalent hybridization in $D_{6h}$ untwisted graphene bilayer, $D_{6}$ twisted graphene bilayer permits equivalent and mixed hybridizations. Furthermore, $D_{6d}$ dodecagonal quasicrystal in $30^{\circ}$-TBG allows for both equivalent and nonequivalent hybridizations [32]. Additionally, the presence of an exchange field and a Rashba spin-orbit coupling have been proposed to induce a nontrivial topological phase in $30^{\circ}$-TBG systems [33].
In this study, we reexamine the electronic properties of $30^{\circ}$-TBGs, with a focus on the spatial distribution of the localized states, using the Haydock-Heine-Kelly (HHK) recursion technique [34, 35, 36, 37]. This method relies on the nearsightedness of electronic matter (see, for instance, Ref. [38]) and, as a real-space approach, unlike the standard techniques [39], eliminates the need for periodic boundary conditions, making it ideal for analyzing strongly disordered, amorphous, and quasicrystalline electronic systems. We demonstrate that the HHK method outperforms the previously proposed approaches for $30^{\circ}$-TBG systems. By developing a geometric model to interpret the LDOS results, we are able to uncover the spatial distribution of the LDOS suppression observed in Ref. [11] and gain insights into the fractal nature of the localized electronic states.
@@ II.1 Lattice structure
Figure 1(a) illustrates the first Stampfli construction method obtained from the grid method [24, 12]. The quasicrystal lattice is composed of equilateral triangles, squares and rhombuses (with $30^{\circ}$ acute angle), arranged with different orientations to maintain a 12-fold rotational symmetry pattern, while lacking translational symmetry. The inflation rule of Stampfli tiles involved rotating the lattice by $15^{\circ}$ around its perfect rotational symmetry center and scaling side lengths by a factor $\sqrt{2+\sqrt{3}}$ [24]. This process generates a new lattice whose vertices coincide with the ones of the original lattice.
@@
Figure 1(b) presents the second construction method as a compound tessellation [30, 24]. The dodecagonal compound tessellation is decomposed into equilateral triangles and squares, as depicted in Fig. 1(c). The center of each dodecagon corresponds to a vertex of a larger, similar dodecagonal element, with side lengths scaled by a factor $2+\sqrt{3}$ [24].
In the next section, we use the properties of the Stampfli construction methods to discuss electronic emergent localization features of $30^{\circ}$-TBG systems. More specifically, we employ the third inflation of Stampfli tiles, seed multiplied by a factor $\left(2+\sqrt{3}\right)^{3/2}$, as used in the first construction method. This ensures that the internal structures of each tile kind are similar [12].
Figure 2: Geometric model: The five different highlighted-color, marked as $r_{1}$, $r_{2}$, $r_{3}$, $c_{1}$, $c_{2}$, indicate an enhanced concentration of carbon atom sites at five different radial distances, using as references (a) the Stampfli tile vertices and (b) the dodecagonal centers obtained by the first and second construction methods, respectively. Red and blue lines correspond to the top and bottom graphene layers, respectively.
Let us introduce a geometric model that identifies the main locations of enhanced concentration of lattice sites. Figure 2 highlights five distinct radial distances denoted as $r_{1}$, $r_{2}$, $r_{3}$, $c_{1}$ and $c_{2}$, where the Stampfli tile vertices serve as central points of emergent geometric site concentration. Central points formed by a single kind of Stampli tiles, coinciding with the main twist axis perpendicular to the plane, exhibit a perfect 12-fold symmetry. In contrast, central points formed by the vertices of different Stampfli tiles, corresponding to carbon atom sites at the same radial distances, are characterized by imperfect rings. Most imperfect rings exhibit an imbalance in emergent state measures, as detailed in Sec. III. The imbalance primarily originates from the distinct symmetry groups of the Stampfli tiles: rhombuses, triangles and squares possess twofold, threefold and fourfold symmetry, respectively, see Fig. 2(a). Near-perfect rings are readily identified by the dodecagonal centers used in the second construction method, red circles in Fig. 1(b), where the center point regions exhibit pronounced 12-fold symmetry and balanced emergent state measures.
@@ II.2 Model Hamiltonian
$V_{pp\pi}(R)=V^{0}_{pp\pi}\exp\left(-\frac{R-a_{0}}{\delta}\right),$ (3)
@@
Variations in the interlayer coupling strength, $V^{0}_{pp\sigma}$, result in a monotonically shift of the energy peaks corresponding to the largest spectral weight of the atomic sites with the strongest interlayer interaction. These states are located at energies far from the Fermi level. On the other hand, variations in the intralayer hopping parameter, $V^{0}_{pp\pi}$, primarily affect the Fermi velocity. The combined influence of both the intralayer and interlayer hopping parameters determines the energy position of the density of states suppression arising from the intersection of two Dirac cones, as discussed in Ref. [11].
@@ II.3 Numerical method
We compute the local spectral function of $30^{\circ}$-TBG quasicristals, described by Eqs. (1)-(5), using the Haydock-Heine-Kelly (HHK) recursion technique [34, 35, 36, 37]. The HHK method is an $O(\mathcal{N})$ approach which consists of a recursive procedure [36] that transforms an arbitrary sparse Hamiltonian matrix in a tridiagonal form. Next, one evaluates the diagonal Green’s function, $G_{ii}^{r}(\epsilon)\equiv G^{r}(\textbf{R}_{i},\textbf{R}_{i},\epsilon)$, by a continued fraction expansion, that is very amenable for numerical calculation. A detailed presentation of the method and its implementation can be found, for instance, in Refs. [36, 37].
@@
As noted in the introduction, the the HHK method is an order $N$ real-space approach that does not require periodic boundary conditions periodic boundary conditions or finite system sizes. Here, we increase the number of considered lattice sites until the LDOS converges, typically with approximately $N\sim 10^{7}$ atomic sites. Although, in principle, the calculation must be repeated for each atomic site, the geometric model introduced above, combined with the self-similarity of the electronic spectral functions for different lattice inflations (see below), indicates that only a finite number of sites need to be considered. This feature is essential for the computational efficiency of the method.
@@ III Results
Using the HHK method we compute the LDOS given by Eq.(6) and confirm that the equivalent sites at identical radial distances share the same local spectral function due to the most relevant symmetry operators, as discussed in Sec. II.1. Consequently, our calculations require the computation of only one twenty-forth of the total number of carbon atoms, within a specified circular region of the $30^{\circ}$-TBG dodecagonal quasicrystal.
@@ III.1 Local electronic structures
In Fig. 4 error bars quantify the standard deviation from the mean value, primarily due to the peaks associated to the valence band singularities that vary along the radial distance. For a given kind of Stampfli tile, the first inflation shows fluctuations of order $10^{-3}$ and $10^{-4}$ along the radial distance. By the third inflation, standard deviations are negligible for all spectral functions, regardless of radial distances. A small standard deviation indicates greater similarity with the first spectral function. Thus, the third inflation is the smallest order at which the spectral functions of the Stampfli unit structures exhibit typical patterns characteristic of the dodecagonal quasicrystal. By accounting for the normalization factor that relates the spectral functions of the different Stampfli unit structures, we observe an excellent agreement within the numerical precision. Our findings suggest that the electronic structure of the system can be calculated using only a single Stampfli unit structure.
@@ III.2 Emergent localization
Figures 6(a)-(e) depict real-space images of emergent localized state of a $30^{\circ}$-TBG system, corresponding to the peaks $\alpha$, $\beta$, $\gamma$, $\tau$, and $\mu$ of DOS singularities in the valence band, respectively. The sizes of the red (blue) circles are proportional to the LDOS at the top (bottom) layer site at which the circle is centered. The peak energies $\alpha$, $\beta$, $\gamma$, $\tau$ and $\mu$ are approximately $-1.163$ eV, $-0.949$ eV, $-0.702$ eV, $-0.842$ eV and $-0.44$ eV, respectively. All emergent states exhibit a non-periodic dodecagonal pattern and can be spatially represented by a quasicrystal lattice model at the third inflation of the Spampfli tiles.
@@
Figures 6(f)-(j) present histograms illustrating the spatial distribution of states corresponding to the energy peaks $\epsilon_{i}$ with $i=\{\alpha,\beta,\gamma,\tau,\mu\}$, respectively. We analyze the variation of emergent localized states along the spacial radial distances, $d=\{r_{1},r_{2},r_{3},c_{1},c_{2}\}$, as indicated in Figure 2, where the vertices of the Stampfli tiles serve as central points. Each histogram shows the average LDOS for the energy peak $\epsilon_{i}$, $\rho(\epsilon_{i},d)=\sum_{j\in d}\rho_{j}(\epsilon_{i})$, where $j$ corresponds to the atomic sites located at a distance $d$. The average LDOS is normalized by its maximum value, $\rho_{\rm max}(\epsilon_{i},d)$.
@@
The black circles in the insets of Figs. 6(a)-(c) and in Figs. 6(d) and 6(e) correspond to a zoom-in view of the structures of the central point regions formed by exclusively by Stampfli rhombuses associated with the energy peaks $\epsilon_{i}$ with $i=\{\alpha,\beta,\gamma,\tau,\mu\}$, respectively. The distribution of the LDOS in each peak, $\rho_{j}(\epsilon_{i})$, exhibits a perfect 12-fold symmetry. In the insets of Figs. 6(a)-(c), the red and green circles identify regions corresponding to dodecagonal centers used in the second construction method and the vertices of different Stampfli tiles used in the first construction method, respectively. While the red circles do not locally preserve 12-fold symmetry, their distribution is close to the latter. However, the green circles deviate significantly from 12-fold symmetry, displaying an imbalance due to the different symmetry groups of the Stampfli tiles.
@@
We observe that the average LDOS at specific radial distances, $\rho(\epsilon_{i},d)$, in regions with strong imbalances resemble those in the regions with perfect 12-fold symmetry.
@@
Figures 6(a) and 6(f) demonstrate that the $\alpha$-peak states are highly localized on the second ring of carbon atoms closest to the center points, at a radial distance $r_{2}$, with $\rho_{\rm max}(\alpha,r_{2})=0.203$ a.u. Interestingly, these second-ring atomic sites can be considered as opposite of dimer sites, due to their significant spacing.
@@
Figures 6(b) and 6(g) reveal a quasi-homogeneous emergent states distribution for the $\beta$-peak, with $\rho_{\rm max}(\beta,c_{2})=0.119$ a.u., indicating less less localization at radial distance $c_{1}$. Figures 6(i) and 6(j) show that the peaks $\tau$ and $\mu$ peaks also exhibit quasi-homogeneous emergent states distributions, with $\rho_{\rm max}(\tau,r_{1})=0.089$ a.u. and $\rho_{\rm max}(\mu,c_{1})=0.062$ a.u., respectively. While the $\tau$-peak states are less localized at $c_{1}$, the $\mu$-peak states are less localized at $r_{2}$ and $c_{2}$.
@@
Figure 6(h) illustrates that the $\gamma$-peak states are highly localized at a radial distance $c_{1}$, coinciding with the dimer sites, with $\rho_{\rm max}(\gamma,c_{1})=0.102$ a.u. In Figure 6(c), the largest localizations are indicated by the purple disks, which correspond to overlapping red and blue disks representing the top and bottom layers, respectively.